# Fibre Inflation Meets Quintessence: Implications of Perturbative Stabilisation

### v1

@@ Abstract
In this paper, we have discovered a new avenue of fibre inflation in perturbative large volume scenario (pLVS) due to the redefinition of the base modulus. pLVS offers a novel regime where large volume of the internal space is guaranteed without the need of non-perturbative effects. In this setup, we study the possibility where a base redefinition allows to assess different versions of fibre inflation whose spectral index aligns with Atacama Cosmology Telescope (ACT) data ACT:2025fju; ACT:2025tim and produces tensor-to-scalar ratio in the range $0.008\lesssim r\lesssim 0.01$ in different setups we have considered. The leading order flat direction - which in our case is the fibre modulus - is lifted with the combinations of string loop corrections, leading order $\alpha^{\prime 3}$$R^{4}$-correction, higher derivative $F^{4}$ corrections as well as our new ingredient redefinition of the modulus. Since recent Dark Energy Spectroscopic Instrument (DESI) results appear to favour a dynamical explanation for late-time acceleration over a simple cosmological constant, exploring quintessence offers a more suitable approach. In this lore, we also examine the quintessence sector to complete our model and account for both early- and late-time cosmic acceleration. In this framework, the poly-instanton correction generates a potential along the axionic directions, and we find that the resulting quintessence behaviour and the subsequent cosmological predictions about dark matter closely resemble the predictions of the original fibre inflation scenario studied earlier.
@@ 1 Introduction
Inflation in its minimal form, is formulated to be driven by a scalar field known as the inflaton that is minimally coupled to gravity. Inflation occurs as the field slowly rolls over its potential, and stops when it nears the minimum of the potential. During inflation, the energy scale can sit within a couple of orders of magnitude below the Planck scale, where Planck-suppressed operators can no longer be ignored. If primordial gravitational waves are large enough to be detectable, the inflaton must traverse a large distances in field space, which pushes any low-energy description beyond its natural radius of convergence unless the ultraviolet (UV) completion dictates the pattern of higher-dimensional corrections. In small-field models, the inflaton traverses a sub-Planckian patch, so observable tensor modes are negligible, and slow-roll then hinges on maintaining a flat potential. Exactly as is the case for inflation, dark energy too, is UV-sensitive: Planck-suppressed operators tend to steepen the potential (the $\eta$ problem), unless protected by a symmetry. String theory, being a UV-complete framework, has the ability to accommodate the necessary ingredients to explain the dynamical nature of gravity along with its quantum aspects. String theory supplies precisely the ingredients which make the study of inflation more robust.
@@
However, one can go further in this story-telling of moduli stabilisation using perturbative corrections. Using the tree-level and one-loop correction at $\alpha^{\prime 3}$, one can stabilise the Kähler moduli Antoniadis:2018hqy; Antoniadis:2019rkh; Leontaris:2022rzj; Leontaris:2023obe. This setup also gives an AdS minima, which can be uplifted via say for example D-terms. A further development was seen in Cicoli:2024bwq, where they stabilised the Kähler modulus directly to a dS minimum using solely perturbative corrections courtesy of moduli redefinitions. Ref. Hai:2025wvs then extended the analysis to multiple Kähler moduli. Interestingly, the tree-level superpotentential, in this perturbative stabilisation scenarios remain unconstrained due to the moduli stabilisation scheme. This brings forth a great sense of confidence when extending to inflationary scenarios. The moduli redefinitions, arising from D7-branes at gauge thresholds Conlon:2009kt; Conlon:2009qa; Conlon:2009xf, offer an interesting possibility in forms of uplifting to dS Cicoli:2024bwq and also in this paper where it pushes fibre inflation to ACT regime.
@@
With moduli stablised, one should then resort to the case of studying both the inflationary regime first. Various inflationary constructions have been proposed in the context of KKLT and LVS (see eg. Baumann:2014nda; Cicoli:2023opf for reviews on these inflationary models). In the perturbative moduli stabilisation scheme, various inflationary scenarios have been developed over the years Bera:2024zsk; Bera:2024sbx; Bera:2024ihl; Chakraborty:2025yms; Leontaris:2025hly; Leontaris:2025xit; Cicoli:2024bwq; Hai:2025wvs. One of the prominent inflationary scenarios is both LVS and pLVS is the Fibre Inflationary scenario [^1] See Cicoli:2008gp; Cicoli:2016xae; Burgess:2016owb; Cicoli:2023njy; Cicoli:2024bxw for fibre inflation in LVS context.. However, as remarked in Ferreira:2025lrd, generic fibre inflation models seem to predict $0.96<n_{s}<0.967$, but according to the current bounds set by ACT ACT:2025tim; ACT:2025fju, the required bounds for $n_{s}$ need to be higher  [^2] A plethora of literatures have tried to address match the ACT data for example in Kallosh:2025rni; Liu:2025qca; Yogesh:2025wak; Peng:2025bws; Yin:2025rrs; Byrnes:2025kit; Wolf:2025ecy; Aoki:2025wld; Gao:2025viy; Zahoor:2025nuq; Ferreira:2025lrd; Mohammadi:2025gbu; Odintsov:2025wai; Q:2025ycf; Zhu:2025twm; Kouniatalis:2025orn; Hai:2025wvs; Dioguardi:2025vci; Yuennan:2025kde; Kallosh:2025gmg; Kallosh:2025ijd; GonzalezQuaglia:2025qem; DOnofrio:2025bol; Oikonomou:2025jmy; Odintsov:2025mqq; Odintsov:2025eiv; Oikonomou:2025xms; Odintsov:2025jky; Oikonomou:2025icu; Odintsov:2025oiy; Nojiri:2025low; Odintsov:2025jfq; Ellis:2025bzi; Yuennan:2025inm; Yuennan:2025tyx; Oikonomou:2025htz; Ellis:2025ieh; Ishiguro:2025tvn; Leontaris:2025hly; Cagan:2025rbc; Maity:2025czp. . In this paper, we try to remedy this situation and push fibre inflation to a more stable ACT-territory.

### v2

@@ Abstract
In this paper, we have discovered a new avenue of fibre inflation in perturbative large volume scenario (pLVS) due to the redefinition of the base modulus. pLVS offers a novel regime where large volume of the internal space is guaranteed without the need of non-perturbative effects. In this setup, we study the possibility where a base redefinition allows to assess different versions of fibre inflation whose spectral index aligns with Atacama Cosmology Telescope (ACT) data and produces tensor-to-scalar ratio in the range $0.008\lesssim r\lesssim 0.01$ in different setups we have considered. The leading order flat direction - which in our case is the fibre modulus - is lifted with the combinations of string loop corrections, leading order $\alpha^{\prime 3}$$R^{4}$-correction, higher derivative $F^{4}$ corrections as well as our new ingredient redefinition of the modulus. Since recent Dark Energy Spectroscopic Instrument (DESI) results appear to favour a dynamical explanation for late-time acceleration over a simple cosmological constant, exploring quintessence offers a more suitable approach. In this lore, we also examine the quintessence sector to complete our model and account for both early- and late-time cosmic acceleration. In this framework, the poly-instanton correction generates a potential along the axionic directions, and we find that the resulting quintessence behaviour and the subsequent cosmological predictions about dark matter closely resemble the predictions of the original fibre inflation scenario studied earlier.
@@ 1 Introduction
Inflation, in its minimal form is formulated to be driven by a scalar field known as the inflaton that is minimally coupled to gravity. Inflation occurs as the field slowly rolls over its potential, and stops when it nears the minimum of the potential. During inflation, the energy scale can sit within a couple of orders of magnitude below the Planck scale, where Planck-suppressed operators can no longer be ignored. If primordial gravitational waves are large enough to be detectable, the inflaton must traverse a large distances in field space, which pushes any low-energy description beyond its natural radius of convergence unless the ultraviolet (UV) completion dictates the pattern of higher-dimensional corrections. In small-field models, the inflaton traverses a sub-Planckian patch, so observable tensor modes are negligible, and slow-roll then hinges on maintaining a flat potential. Exactly as is the case for inflation, dark energy too, is UV-sensitive: Planck-suppressed operators tend to steepen the potential (the $\eta$ problem), unless protected by a symmetry. String theory, being a UV-complete framework, has the ability to accommodate the necessary ingredients to explain the dynamical nature of gravity along with its quantum aspects. String theory supplies precisely the ingredients which make the study of inflation more robust.
@@
However, one can go further in this story-telling of moduli stabilisation using perturbative corrections. Using the tree-level and one-loop correction at $\alpha^{\prime 3}$, one can stabilise the Kähler moduli Antoniadis:2018hqy; Antoniadis:2019rkh; Leontaris:2022rzj; Leontaris:2023obe; Basiouris:2020jgp; Basiouris:2021sdf; Basiouris:2025yir. This setup also gives an AdS minima, which can be uplifted via say for example D-terms. A further development was seen in Cicoli:2024bwq, where they stabilised the Kähler modulus directly to a dS minimum using solely perturbative corrections courtesy of moduli redefinitions. Ref. Hai:2025wvs then extended the analysis to multiple Kähler moduli. Interestingly, the tree-level superpotentential, in this perturbative stabilisation scenarios remain unconstrained due to the moduli stabilisation scheme. This brings forth a great sense of confidence when extending to inflationary scenarios. The moduli redefinitions, arising from D7-branes at gauge thresholds Conlon:2009kt; Conlon:2009qa; Conlon:2009xf, offer an interesting possibility in forms of uplifting to dS Cicoli:2024bwq and also in this paper where it pushes fibre inflation to ACT regime.
@@
With moduli stablised, one should then resort to the case of studying both the inflationary regime first. Various inflationary constructions have been proposed in the context of KKLT and LVS (see eg. Baumann:2014nda; Cicoli:2023opf for reviews on these inflationary models). In the perturbative moduli stabilisation scheme, various inflationary scenarios have been developed over the years Bera:2024zsk; Bera:2024sbx; Bera:2024ihl; Chakraborty:2025yms; Leontaris:2025hly; Leontaris:2025xit; Cicoli:2024bwq; Hai:2025wvs. One of the prominent inflationary scenarios is both LVS and pLVS is the Fibre Inflationary scenario [^1] See Cicoli:2008gp; Cicoli:2016xae; Burgess:2016owb; Bhattacharya:2020gnk; Cicoli:2023njy; Cicoli:2024bxw for fibre inflation in LVS context.. However, as remarked in Ferreira:2025lrd, generic fibre inflation models seem to predict $0.96<n_{s}<0.967$, but according to the current bounds set by ACT ACT:2025tim; ACT:2025fju, the required bounds for $n_{s}$ need to be higher  [^2] A plethora of literatures have tried to address matching ACT data for example in Kallosh:2025rni; Liu:2025qca; Yogesh:2025wak; Peng:2025bws; Yin:2025rrs; Byrnes:2025kit; Wolf:2025ecy; Aoki:2025wld; Gao:2025viy; Zahoor:2025nuq; Ferreira:2025lrd; Mohammadi:2025gbu; Odintsov:2025wai; Q:2025ycf; Zhu:2025twm; Kouniatalis:2025orn; Hai:2025wvs; Dioguardi:2025vci; Yuennan:2025kde; Kallosh:2025gmg; Kallosh:2025ijd; GonzalezQuaglia:2025qem; DOnofrio:2025bol; Oikonomou:2025jmy; Odintsov:2025mqq; Odintsov:2025eiv; Oikonomou:2025xms; Odintsov:2025jky; Oikonomou:2025icu; Odintsov:2025oiy; Nojiri:2025low; Odintsov:2025jfq; Ellis:2025bzi; Yuennan:2025inm; Yuennan:2025tyx; Oikonomou:2025htz; Ellis:2025ieh; Ishiguro:2025tvn; Leontaris:2025hly; Cagan:2025rbc; Maity:2025czp; Gialamas:2025kef; Gialamas:2025ofz. . In this paper, we try to remedy this situation and push fibre inflation to a more stable ACT-territory.